OpenStudy (anonymous):

please help how do i express 5/(3+sqrt 2) in the form of a+b sqrt 2 ?

OpenStudy (anonymous):

5/(3+sqrt 2) = 5*(3- sqrt 2) /( (3+sqrt 2)* (3-sqrt 2) ) =5*(3- sqrt 2) /(3^2 - (sqrt2)^2) = 5*(3- sqrt 2) /(9-2) = 5*(3- sqrt 2) /7 =15/7 + ( -5/7) *sqrt 2

OpenStudy (anonymous):

thank you